OpenStudy (shelby1290):

Solve using the quadratic formula: x^2+7x=24

MsBrains (ms-brains):

If you list the whole problem. I can help better

OpenStudy (shelby1290):

@Ms-Brains what do you mean by "listing the whole problem" ?

OpenStudy (anonymous):

@shelby1290 I can help :D

MsBrains (ms-brains):

Take a snip of it, and attach the link. :)

OpenStudy (shelby1290):

@AskMeLotss thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (shelby1290):

@Ms-Brains this IS the math problem. It's an application question. What more can I possibly say?

OpenStudy (jchick):

Step 1 : Simplify x2 - 7x Equation at the end of step 1 : (x2 - 7x) - 24 = 0

MsBrains (ms-brains):

The list of choices.. But never mind. cx You already git help. :3

OpenStudy (anonymous):

what are the answer choices? if you dont have any, i can still help also :)

OpenStudy (shelby1290):

It isn't a multiple choice question though.....

OpenStudy (jchick):

Step 2 : Simplify x2-7x - 24 Trying to factor by splitting the middle term

OpenStudy (jchick):

Factoring x2-7x-24 The first term is, x2 its coefficient is 1 . The middle term is, -7x its coefficient is -7 . The last term, "the constant", is -24

OpenStudy (jchick):

Step-1 : Multiply the coefficient of the first term by the constant 1 • -24 = -24

OpenStudy (shelby1290):

It's an open response/application question. There aren't any choices.

OpenStudy (jchick):

Step-2 : Find two factors of -24 whose sum equals the coefficient of the middle term, which is -7

OpenStudy (jchick):

-24 + 1 = -23 -12 + 2 = -10 -8 + 3 = -5 -6 + 4 = -2 -4 + 6 = 2 -3 + 8 = 5 -2 + 12 = 10 -1 + 24 = 23 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored

OpenStudy (jchick):

Equation at the end of step 2 : x2 - 7x - 24 = 0 Step 3 : Solve x2-7x-24 = 0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Solve+using+the+quadratic+formula+x%5E2%2B7x%3D24 click this and it will help you get your answer, it HAS the answer

OpenStudy (jchick):

Solving x2-7x-24 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 1 B = -7 C = -24 Accordingly, B2 - 4AC = 49 - (-96) = 145 Applying the quadratic formula : 7 ± √ 145 x = ————— 2 √ 145 , rounded to 4 decimal digits, is 12.0416 So now we are looking at: x = ( 7 ± 12.042 ) / 2 Two real solutions: x =(7+√145)/2= 9.521 or: x =(7-√145)/2=-2.521

OpenStudy (shelby1290):

@jchick thank you so much

OpenStudy (jchick):

No problem!